OpenStudy (anonymous):

cos(sin^-1(3/7))=_________ A. 3/2√10 B. 7/2√10 C. 2√10/7 D. Undefined

OpenStudy (epoweritheta):

let sin-1 3/7=x => sinx =3/7 => cosx=root(40)/7 =>cox(sin-1 3/7)=2√10/7

OpenStudy (anonymous):

@epoweritheta can you help me with another one?

OpenStudy (epoweritheta):

ya...

OpenStudy (anonymous):

\[\tan(\sin ^{-1}(-1)\] A. 1 B. 0 C. -1 D. Undefined

OpenStudy (epoweritheta):

if sinx=-1 ,what is x ??? in -pi tp pi ??

OpenStudy (anonymous):

What is tp?

OpenStudy (anonymous):

@epoweritheta

OpenStudy (epoweritheta):

sorry to

OpenStudy (epoweritheta):

not tp

OpenStudy (anonymous):

Ok I am confused